OpenStudy (anonymous):

which is true? 1.4 < 1 1/5 1 3/20 > 1.4 2 3/5 >2.8 Orrrrrrr 2.8 < 2 17/20 (Note< I have put this up like....5 mins ago but I got up and my sister thought it would be funny to shut my laptop down... Can you pleaseee help me!

OpenStudy (anonymous):

Do you know how to solve this?

OpenStudy (hawkfalcon):

... I don't even... I could do this in first grade;o

jimthompson5910 (jim_thompson5910):

Convert each number to decimal form Ex: \[\Large 1\frac{1}{5} = 1+\frac{1}{5}\] \[\Large 1\frac{1}{5} = 1+0.2\] \[\Large 1\frac{1}{5} = 1.2\]

jimthompson5910 (jim_thompson5910):

So \[\Large 1.4 < 1\frac{1}{5}\] becomes \[\Large 1.4 < 1.2\] which is false

OpenStudy (anonymous):

:)